OpenStudy (anonymous):

Find three fractions equivalent to each of the following: 1/2 4/5 2/3 5/6 7/8 9/10

OpenStudy (asnaseer):

use the fact that a fractions value remains the same if you multiply its numerator and denominator by the same number. so, for example:\[\frac{1}{5}=\frac{1}{5}*\frac{3}{3}=\frac{3}{15}\]

OpenStudy (anonymous):

Yep ^ He's right